eviant:

If f(x) = 4x – 9, what is the equation for f–1(x)?

eviant:

@dude

dude:

Flip x and y and solve for y \(x=4y-9\)

eviant:

Add 9 to both sides:x+9=4y Divide both sides by 4:\[\frac{ x+9 }{ 4 }\]

eviant:

y=\[\frac{ x+9 }{4 }\]

eviant:

@dude ?

dude:

Right